OpenStudy (anonymous):

A fireworks company has two types of rockets called Zinger 1 and Zinger 2. The polynomial gives the height in feet of Zinger 1 at t seconds after launch. The polynomial: -16t2 + 165t gives the height of Zinger 2 at t seconds after launch. If the rockets are launched at the same time and both explode 6 seconds after launch, how much higher is Zinger 2 than Zinger 1 when they explode? A. 414 ft B. 990 ft C. 90 ft D. 324 ft

OpenStudy (anonymous):

PLEASE HELP!!!!

OpenStudy (anonymous):

what is the polynomial representing 1

OpenStudy (anonymous):

16t^2+150t

OpenStudy (anonymous):

there is a -ve sign before Z2's polynomial?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

there is no -ve sign before t^2 for Z1?

OpenStudy (anonymous):

A fireworks company has two types of rockets called Zinger 1 and Zinger 2. The polynomial -16t^2+150t gives the height in feet of Zinger 1 at t seconds after launch. The polynomial: -16t2 + 165t gives the height of Zinger 2 at t seconds after launch. If the rockets are launched at the same time and both explode 6 seconds after launch, how much higher is Zinger 2 than Zinger 1 when they explode?

OpenStudy (anonymous):

put t=6 distance that Z1 travels = -16(6)^2+150(6) distance that Z2 travels = -16(6)^2 + 165(6) distance between them = -16(6)^2 + 165(6) - (-16(6)^2+150(6)) = 90

OpenStudy (anonymous):

A. 414 ft B. 990 ft C. 90 ft D. 324 ft these r the answers

OpenStudy (anonymous):

cool, Thank you so much :P

OpenStudy (anonymous):

you are welcome

OpenStudy (anonymous):

:D